OpenStudy (anonymous):

if f(x)=2x+8 and g(x)=x^4, what is (gof)(-3)

OpenStudy (anonymous):

the o being the little circle thing

Parth (parthkohli):

\(g \circ f (x) = g(f(x)) \)

Parth (parthkohli):

\(g(2x + 8) \Longrightarrow (2x + 8)^4\)

Parth (parthkohli):

\(g(f(-3))\)

OpenStudy (lgbasallote):

first find f(-3) \[f(-3) = 2(-3) + 8 \implies -6 + 8 \implies 2\] now find g(2) \[g(f(-3)) = g(2) = (2)^4 \implies 2 \times 2\times 2 \times 2 \implies 16\]

Parth (parthkohli):

That'd look like \(g(f(x)) = (2x + 8)^4 \Longrightarrow (2(-3) + 8)^4\)

OpenStudy (anonymous):

thanks for the quick reply parth! Good answer lg

Parth (parthkohli):

That's right. It's an answer.